# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

### v1

@@ 1 Introduction
In the proofs that follow we rely on a series of small lemmas which are stated and proven in the Appendix. We also reference Equation (20) to emphasize the technique used to rewrite an $L^{2}$ norm into scalar factors.
@@
    Notice that there always exists a constant $K>0$, independent of both $\lambda_{\ell}$ and $r>1$, such that $|\gamma_{k\ell}(r)^{2}|=|\lambda_{\ell}+\mu_{k}/r^{2}|\leq K(1+\mu_{k})$. Then, by Equation (20), we have that
@@
    Let $r_{1},r_{2}\geq 1$ be sufficiently close so that $r_{1}\leq 2r_{2}$. Then
    $\|A(r_{1})^{1/2}u-A(r_{2})^{1/2}u\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
@@
    We claim that the operator $\partial_{r}\!\left[A(r)^{1/2}\right]$, defined via the symbol $\partial_{r}\gamma_{k\ell}(r)=\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)}\right)$, is the Fréchet derivative of $A(r)^{1/2}$. To show this, fix $r\geq 1$ and let $r^{\prime}\geq 1$ be sufficiently close to $r$ so that $r^{\prime}/2\leq r<2r^{\prime}$. Then
@@
    $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)-\gamma_{k\ell}(r^{\prime})-\partial_{r}\gamma_{k\ell}(r)(r-r^{\prime})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
@@
    $\lim_{r^{\prime}\to r}\frac{\|A(r)^{1/2}-A(r^{\prime})^{1/2}-\partial_{r}A(r)^{1/2}(r-r^{\prime})\|_{L^{2}}}{|r-r^{\prime}|}\leq\lim_{r^{\prime}\to r}\frac{C^{\prime}|r-r^{\prime}|^{2}}{|r-r^{\prime}|}=0.$
@@
    Let $r_{1}\geq r_{2}\geq 1$ be sufficiently close so that $r_{2}/2\leq r_{1}\leq 2r_{2}$. Then
    $\|\partial_{r}A(r_{1})^{1/2}u-\partial_{r}A(r_{2})^{1/2}u\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|^{2}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}$ by Eq (20)
    $\leq\tilde{C}\left||r_{1}-r_{2}|+|r_{1}^{3}-r_{2}^{3}|\right|^{2}\sum_{k,j}(1+\mu_{k})\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}\quad$ by Lemma A.6
@@
    First note that the symbol of $A(r)^{-1/2}$ is given by $\gamma_{k\ell}(r)^{-1}$ and that consequently the symbol of the product $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is given by $\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right)$. We therefore obtain a bound on the $L^{2}$ operator norm of $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ by bounding the supremum of the symbols over all values of $k$ and $\ell$. This bound, given by
@@
    exists with $C$ independent of $r$ by hypothesis (H2.). Consequently, by Equation (20),
    $\left\|\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}u\right\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left|\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right|\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}\leq\tilde{C}r^{-1}\sum_{kj}\|u_{kj}\|_{\mathbb{C}^{d}}^{2}=\tilde{C}r^{-1}\|u\|_{L^{2}}^{2}$
    with $\tilde{C}$ also independent of $r$, and the operator norm on $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is $O(r^{-1})$.
To prove (ii), we first prove that $\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$ is bounded above by a constant times $\|\mathbf{u}\|_{X_{r}}^{2}$:
$\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left\{\left(|\lambda_{\ell}|+\mu_{k}/r^{2}\right)\right\}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}+\|v\|_{L^{2}}^{2}$ by Eq (20)
[…]
@@
The proof of Lemma 2.5 uses the same spectral decomposition ideas to construct an exponential dichotomy for (17) from the hyperbolicity of $A_{\infty}$ given by (H2).
@@
where by the spectral structure of $V_{\infty}$, we can further decompose (21) according to
$\mathrm{exp}\left(\pm\int_{t}^{r}A_{k}(s)^{1/2}ds\right)=R\begin{pmatrix}\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}.$ (22)
The continuity and differentiability of (21) follows from the continuity and differentiability of both $A(r)^{1/2}$ and each $\gamma_{k\ell}(r)$ (Lemmas 2.4, A.4, A.5, and A.6). By taking the derivatives of $\tilde{\Phi}^{\mathrm{s}}(r,t)\mathbf{w}$ and $\tilde{\Phi}^{\mathrm{u}}(r,t)\mathbf{w}$ with respect to $r$ one can directly confirm that for each $t\in\mathbb{R}$ and $\mathbf{w}\in(L^{2}(S^{n-1})\times L^{2}(S^{n-1}))$ these are both solutions to (17). The invariance conditions of Definition 2.1 are also clearly satisfied by the evolutionary operators and their corresponding projections, $\tilde{P}^{\mathrm{s}}:=\tilde{\Phi}^{\mathrm{s}}(r,r)=\left(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\right)$ and $\tilde{P}^{\mathrm{u}}:=\tilde{\Phi}^{\mathrm{u}}(r,r)=\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)$.
It remains to be shown that the exponential decay estimates hold for $\tilde{\eta}:=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]$. These decay estimates follow from the fact that for all $k\in\{0,1,2,\ldots\}$ and $r>0$, $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]\geq\mathrm{Re}[\lambda_{\ell}^{1/2}]$ (see Lemma A.8). We therefore obtain:
$\left\|\tilde{\Phi}^{s}(r,t)\mathbf{w}\right\|_{L^{2}\times L^{2}}^{2}$ $=\sum_{k,j}\left\|\mathrm{exp}\left(-\int_{t}^{r}A_{k}(s)^{1/2}ds\right)w^{-}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
@@
$\leq\sum_{k,j}\|R\|^{2}\|R^{-1}\|^{2}\left(\max_{1\leq\ell\leq d}\left[e^{-\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right]^{2}\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}C\left(e^{-\min\limits_{1\leq\ell\leq d}2\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
@@
$\left\|\tilde{\Phi}^{u}(r,t)\mathbf{w}\right\|_{L^{2}\times L^{2}}^{2}$ $=\sum_{k,j}\left\|\mathrm{exp}\left(\int_{t}^{r}A_{k}(s)^{1/2}ds\right)w^{+}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$=\sum_{k,j}\left\|R\begin{pmatrix}\mathrm{exp}\left(-\int_{r}^{t}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(-\int_{r}^{t}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}w^{+}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\|R\|^{2}\left(\max_{1\leq\ell\leq d}\left|\mathrm{exp}\left(-\int_{r}^{t}\gamma_{k\ell}(s)ds\right)\right|^{2}\right)\|R^{-1}\|^{2}\|w^{+}_{kj}\|^{2}_{\mathbb{C}^{d}}$
$\leq\sum_{k,j}\|R\|^{2}\|R^{-1}\|^{2}\left(\max_{1\leq\ell\leq d}\left[e^{-\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](t-r)}\right]^{2}\right)\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}C\left(e^{-\min\limits_{1\leq\ell\leq d}2\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](t-r)}\right)\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}Ce^{-2\tilde{\eta}(t-r)}\|w^{+}_{kj}\|_{\mathbb{C}^{d}}^{2}=Ce^{2\tilde{\eta}(r-t)}\|w^{+}\|_{L^{2}}^{2}\leq Ce^{2\tilde{\eta}(r-t)}\|\mathbf{w}\|_{L^{2}\times L^{2}}^{2},$ $r\leq t$
@@
From (23) we can directly apply Proposition 2.3 to get an exponential dichotomy $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$ for (16) in $L^{2}\!\!\times\!L^{2}$ for $r>1$ again with the same decay rate $\tilde{\eta}$, noting that the perturbed part of (16) is $\mathcal{O}(1/r)$.
@@
Notice that the transformation $(u,v)\mapsto(A(r)^{1/2}u,v)$ and can be written as
@@
Therefore, (16) is equivalent to $\mathbf{\tilde{u}}_{r}=\big[T(r)\tilde{\mathcal{A}}(r)T(r)^{-1}+\partial_{r}\left[T(r)\right]T(r)^{-1}\big]\mathbf{\tilde{u}}$. A similar $r$-dependent conjugation by $T$ will transform $\hat{\Phi}^{\mathrm{s}}_{*}(r,t)$ and $\hat{\Phi}^{\mathrm{u}}_{*}(r,t)$ from a family of solution operators for (16) in $L^{2}\times L^{2}$ to a family of solution operators for (9) in $X^{r}$.
@@
$\Phi^{\mathrm{u}}_{*}(r,t)$ $:=T(r)^{-1}\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t):X_{t}\to X_{r},\qquad r\leq t$ (24)
@@
Though the isomorphism $T(r)$ is not an isometry, by Lemma 2.4 it does respect the norm up to linear constants. Therefore the family $\Phi_{*}^{\mathrm{s}/\mathrm{u}}$ inherits the same exponential decay properties in $X_{r}$ as the exponential dichotomy $\hat{\Phi}_{*}^{\mathrm{s}/\mathrm{u}}$ has in $L^{2}\!\!\times\!L^{2}$, including the same decay rate $\tilde{\eta}$. For any $\mathbf{u}\in X_{t}$,
$\left\|\Phi_{*}^{\mathrm{s}}(r,t)\mathbf{u}\right\|_{X_{r}}^{2}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{s}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq C_{2}Ke^{-\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq\frac{C_{2}K}{C_{1}}e^{-\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}}^{2},$ $r\geq t$
and
$\left\|\Phi_{*}^{\mathrm{u}}(r,t)\mathbf{u}\right\|_{X_{r}}^{2}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq C_{2}Ke^{\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}^{2}\leq\frac{C_{2}K}{C_{1}}e^{\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}}^{2},$ $r\leq t.$
@@
Theorem 2.1, together with the results of [1], allow one to fully characterize the bounded behavior of spatial evolutionary systems that arise from elliptic PDEs of the form $\Delta u+V(x)u=0$ where $V(x)=V(r,\cdot)$ decays sufficiently as $r$ goes to infinity. In fact, exponential dichotomies provide a construction of bounded solutions to nonlinearly perturbed equations via a Duhamel formula. This gives rise to a constructive approach to proving the existence to localized stationary solutions of pattern-forming PDEs that lack any kind of rotational or dihedral symmetry.
To see the broad applicability of Theorem 2.1, notice that any stationary reaction diffusion model of the form $0=\Delta u+f(u)$, $u\in\mathbb{C}^{d}$, becomes $0=\Delta u+Df(q(x))u$ when linearized about a stationary solution $q(x)$. Notice also that this matrix $Df(q(x))$ will certainly decay to the constant coefficient matrix $Df(0)$ if the solution $q(x)$ is localized. Therefore the linearization of such a reaction diffusion stationary problem on $\mathbb{R}^{n}$ about any localized solution $q(x)$ has exponential dichotomies in both the core field and the far field provided that $Df(0)$ has eigenvalues away from the negative real line. We could then perturb the reaction diffusion model, including a small spatially inhomogeneous nonlinearity,
$0=\Delta u+f(u)+\mu h(x,u,\mu)=\Delta u+Df(q(x))u+\big[f(u)-Df(q(x))u+\mu h(x,u,\mu)\big]$ (25)
and use the core field and far field dichotomies for the linearized problem to construct bounded solutions to (25) in the core and far field, respectively. Showing that those two solutions connect to form a single bounded solution on the whole domain requires more work.
This type of construction was demonstrated in the channel setting, $(x,y)\in\mathbb{R}\times\Omega$, by Peterhof, Sandstede, and Scheel in [11] for a spatial evolutionary equation of the form:
@@
where $A$ is an autonomous, linear (differential) operator on $L^{2}(\Omega)$ and $H$ is periodic in $x$. They assume that a homoclinic solution $q(x)$ exists for $\mu=0$ in (26), and that the linearization about $q$ has an exponential dichotomy on $\mathbb{R}^{+}$ and on $\mathbb{R}^{-}$. They proceed by constructing a bounded solution on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$ respectively and proving via a Melnikov argument that the stable space on $\mathbb{R}^{+}$ and the unstable space on $\mathbb{R}^{-}$ intersect in such a way that a unique bounded solution exists on the whole domain $\mathbb{R}$ when $\mu>0$ is small.
Ongoing and future work aims to extend this Melnikov argument to our higher dimensional context with the primary challenge being the higher dimensional kernel that comes from having not just one translational symmetry but $n$ translation symmetries, one for each unbounded direction in state space. Once such an argument is completed, the existence of exponential dichotomies in the setting described above would allow us to take previously understood radially symmetric localized solutions and perturb them in non-radially symmetric ways to prove the existence of arbitrary localized stationary patterns.
@@
For all $k>0$, $\mu_{k}^{1/2}\leq(1+\mu_{k})^{1/2}$. Furthermore, $\sum_{k,j}\mu_{k}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}\leq\|u\|_{H^{1}}^{2}$.

### v2

@@ 1 Introduction
In the proofs that follow we rely on a series of lemmas to bound scalar terms which are stated and proven in the Appendix. We also rely repeatedly on the spectral decomposition and diagonalization of $A(r)^{1/2}$ to decompose the $L^{2}(S^{n-1})$ norm as in (18).
@@
    Notice that there always exists a constant $K>0$, independent of both $\lambda_{\ell}$ and $r>1$, such that $|\gamma_{k\ell}(r)^{2}|=|\lambda_{\ell}+\mu_{k}/r^{2}|\leq K(1+\mu_{k})$. Then, by (18), we have
@@
    Let $r_{1},r_{2}\geq 1$ be sufficiently close with $r_{1}\leq 2r_{2}$. Then, again by our Fourier-like decomposition of the $L^{2}(S^{n-1})$ norm, as in (18), we have
    $\|A(r_{1})^{1/2}u-A(r_{2})^{1/2}u\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r_{1})-\gamma_{k\ell}(r_{2})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$
@@
    We claim that the operator $\partial_{r}\!\left[A(r)^{1/2}\right]$, defined via the symbol $\partial_{r}\gamma_{k\ell}(r)=\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)}\right)$, is the Fréchet derivative of $A(r)^{1/2}$. To show this, fix $r\geq 1$ and let $r^{\prime}\geq 1$ be sufficiently close to $r$ so that $r/2\leq r^{\prime}\leq 2r$. Then, again by our Fourier-like decomposition, we have
@@
    $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)-\gamma_{k\ell}(r^{\prime})-\partial_{r}\gamma_{k\ell}(r)(r-r^{\prime})|^{2}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}$
@@
    $\lim_{r^{\prime}\to r}\frac{\|A(r)^{1/2}-A(r^{\prime})^{1/2}-\partial_{r}A(r)^{1/2}(r-r^{\prime})\|_{L(L^{2})}}{|r-r^{\prime}|}\leq\lim_{r^{\prime}\to r}\frac{C^{\prime}|r-r^{\prime}|^{2}}{|r-r^{\prime}|}=0.$
@@
    Let $r_{1}\geq r_{2}\geq 1$ be sufficiently close so that $r_{2}/2\leq r_{1}\leq 2r_{2}$. Then, we have
    $\|\partial_{r}A(r_{1})^{1/2}u-\partial_{r}$ $A(r_{2})^{1/2}u\|_{L^{2}}^{2}\leq\sum_{k,j}C\max_{1\leq\ell\leq d}|\partial_{r}\gamma_{k\ell}(r_{1})-\partial_{r}\gamma_{k\ell}(r_{2})|^{2}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}$
[…]
@@
    First note that the symbol of $A(r)^{-1/2}$ is given by $\gamma_{k\ell}(r)^{-1}$ and that, consequently, the symbol of the product $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is given by $\left(\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right)$. We therefore obtain a bound on the $L^{2}$ operator norm of $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ by bounding the supremum of the symbols over all values of $k$ and $\ell$. We find
@@
    for some $C>0$ independent of $r$ by Hypothesis (H2). Consequently, by our Fourier-like decomposition of the $L^{2}(S^{n-1})$ norm, we have
    $\left\|\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}u\right\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left|\frac{-\mu_{k}/r^{3}}{\gamma_{k\ell}(r)^{2}}\right|\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}$
    $\leq\tilde{C}r^{-1}\sum_{kj}\|u_{kj}\|_{\mathbb{C}^{d}}^{2}=\tilde{C}r^{-1}\|u\|_{L^{2}}^{2},$
    with $\tilde{C}$ also independent of $r$, so the operator norm on $\partial_{r}\!\left[A(r)^{1/2}\right]\!A(r)^{-1/2}$ is $O(r^{-1})$.
To prove (ii), we first apply (18) and estimate that
$\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}$ $\leq\sum_{k,j}C\max_{1\leq\ell\leq d}\left\{\left(|\lambda_{\ell}|+\mu_{k}/r^{2}\right)\right\}\|u_{k,j}\|^{2}_{\mathbb{C}^{d}}+\|v\|_{L^{2}}^{2}$
$\leq\tilde{C}\left(\max_{1\leq\ell\leq d}\left\{|\lambda_{\ell}|\right\}\|u\|_{L^{2}}^{2}+\frac{1}{r^{2}}\|u\|_{H^{1}}^{2}\right)+\|v\|_{L^{2}}^{2}\leq\tilde{C}\|\mathbf{u}\|_{X_{r}}^{2}.$
@@
The proof of Lemma 2.4 explicitly constructs an exponential dichotomy for (17) from the hyperbolicity of $\mathcal{A}_{\infty}$ by exploiting the spectral decomposition and diagonalization of $A(r)^{1/2}$.
@@
where by the diagonalization of $V_{\infty}$, we can further decompose (19) according to
$\mathrm{exp}\left(\pm\int_{t}^{r}A_{k}(s)^{1/2}ds\right)=R\begin{pmatrix}\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{k1}(s)^{1/2}ds\right)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\mathrm{exp}\left(\pm\int_{t}^{r}\gamma_{kd}(s)^{1/2}ds\right)\end{pmatrix}R^{-1}.$
The continuity and differentiability of (19) follows from the continuity and differentiability of both $A(r)^{1/2}$ and each $\gamma_{k\ell}(r)$ (Lemmas 2.3, A.4, A.5, and A.6). By taking the derivatives of $\tilde{\Phi}^{\mathrm{s}}(r,t)\mathbf{w}$ and $\tilde{\Phi}^{\mathrm{u}}(r,t)\mathbf{w}$ with respect to $r$ one can directly confirm that for each $t\in\mathbb{R}$ and $\mathbf{w}\in(L^{2}(S^{n-1})\times L^{2}(S^{n-1}))$ these are both solutions to (17). The invariance conditions required by exponential dichotomies (as in Definition 2.1) are also clearly satisfied by the evolutionary operators and their corresponding projections, $\tilde{P}^{\mathrm{s}}:=\tilde{\Phi}^{\mathrm{s}}(r,r)=\left(\begin{smallmatrix}0&0\\ 0&1\end{smallmatrix}\right)$ and $\tilde{P}^{\mathrm{u}}:=\tilde{\Phi}^{\mathrm{u}}(r,r)=\left(\begin{smallmatrix}1&0\\ 0&0\end{smallmatrix}\right)$.
It remains to be shown that the exponential decay estimates hold for $\tilde{\eta}:=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]$. These decay estimates follow from the fact that for all $k\in\{0,1,2,\ldots\}$ and $r>0$, the following bound holds: $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]\geq\mathrm{Re}[\lambda_{\ell}^{1/2}]$ (see Lemma A.8). We therefore obtain:
$\left\|\tilde{\Phi}^{\mathrm{s}}(r,t)\mathbf{w}\right\|_{L^{2}\times L^{2}}^{2}$ $=\sum_{k,j}\left\|\mathrm{exp}\left(-\int_{t}^{r}A_{k}(s)^{1/2}ds\right)w^{-}_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
@@
$\leq\sum_{k,j}\|R\|^{2}\|R^{-1}\|^{2}\left(\max_{1\leq\ell\leq d}\left[\mathrm{exp}\left({-\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right)\right]^{2}\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}C\left(\mathrm{exp}\left({-\min\limits_{1\leq\ell\leq d}2\mathrm{Re}\big[\lambda_{\ell}^{1/2}\big](r-t)}\right)\right)\|w^{-}_{kj}\|_{\mathbb{C}^{d}}^{2}$
@@
A nearly identical calculation shows the analogous statement for $\tilde{\Phi}^{\mathrm{u}}(r,t)$ on $r\leq t$. ∎
@@
From (20) we can directly apply the robustness of exponential dichotomies (Theorem 1 in [17]) to get an exponential dichotomy $\hat{\Phi}^{\mathrm{s}/\mathrm{u}}_{*}$ for (16) in $L^{2}\times L^{2}$ for $r>1$ again with the same decay rate $\tilde{\eta}$, noting that the perturbed part of (16) is $\mathcal{O}(1/r)$.
@@
Notice that by Lemma 2.3 (ii) the transformation $(u,v)\mapsto(A(r)^{1/2}u,v)$ and can be written as
@@
Therefore, (16) is equivalent to $\mathbf{\tilde{u}}_{r}=\big[T(r)\tilde{\mathcal{A}}(r)T(r)^{-1}+\partial_{r}\left[T(r)\right]T(r)^{-1}\big]\mathbf{\tilde{u}}$.
A similar $r$-dependent conjugation by $T$ will transform $\hat{\Phi}^{\mathrm{s}}_{*}(r,t)$ and $\hat{\Phi}^{\mathrm{u}}_{*}(r,t)$ from a family of solution operators for (16) in $L^{2}\times L^{2}$ to a family of solution operators for (14) in $X^{r}$:
@@
$\Phi^{\mathrm{u}}_{*}(r,t)$ $:=T(r)^{-1}\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t):X_{t}\to X_{r},\qquad r\leq t.$ (21)
@@
Though the isomorphism $T(r)$ is not an isometry, by Lemma 2.3 it does respect the norm up to constants independent of $r$. Therefore the family $\Phi_{*}^{\mathrm{s}/\mathrm{u}}$ inherits the same exponential decay properties in $X_{r}$ as the exponential dichotomy $\hat{\Phi}_{*}^{\mathrm{s}/\mathrm{u}}$ has in $L^{2}\!\!\times\!L^{2}$, including the same decay rate $\tilde{\eta}$. For any $\mathbf{u}\in X_{t}$,
$\left\|\Phi_{*}^{\mathrm{s}}(r,t)\mathbf{u}\right\|_{X_{r}}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{s}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}\leq C_{2}Ke^{-\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}\leq\frac{C_{2}K}{C_{1}}e^{-\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}},$ $r\geq t,$
$\left\|\Phi_{*}^{\mathrm{u}}(r,t)\mathbf{u}\right\|_{X_{r}}$ $\leq C_{2}\left\|\hat{\Phi}_{*}^{\mathrm{u}}(r,t)T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}\leq C_{2}Ke^{\tilde{\eta}(r-t)}\left\|T(t)\mathbf{u}\right\|_{L^{2}\!\times L^{2}}\leq\frac{C_{2}K}{C_{1}}e^{\tilde{\eta}(r-t)}\left\|\mathbf{u}\right\|_{X_{t}},$ $r\leq t.$
@@
Theorem 2.1, together with the results of [3], aid in the characterization of the bounded behavior of spatial evolutionary systems that arise from elliptic PDEs of the form $\Delta u+V(x)u=0$ where $V(x)=V(r,\cdot)$ decays to a constant potential sufficiently fast as $r$ goes to infinity. In fact, exponential dichotomies provide a construction of bounded solutions to nonlinearly perturbed equations via a Duhamel formula. This gives rise to a constructive approach to proving the existence of localized stationary solutions of pattern-forming PDEs that lack any kind of constraining symmetry. Radially symmetric localized solutions were thoroughly characterized by [24]. Recent work in proving the existence of localized patterns with dihedral symmetries (e.g. hexagonal patterns) utilized a combination of analytic and numerical techniques; see [5] for a comprehensive review. In [6], computer-assisted proof techniques via Newton-Kantorovich methods have recently proven successful for rigorously verifying the existence of localized stationary solutions to Swift-Hohenburg without restricting to specific symmetries. However, with our contribution of far-field exponential dichotomies in the setting of unbounded multidimensional spatial domains, we hope to open up the tools of spatial dynamics as set forward by [17] and [21] to this new setting.
To see the broad applicability of Theorem 2.1, notice that any stationary reaction diffusion model of the form $0=\Delta u+f(u)$, $u\in\mathbb{C}^{d}$, becomes $0=\Delta u+Df(q(x))u$ when linearized about a stationary solution $q(x)$. Notice also that this matrix $Df(q(x))$ will certainly decay to the constant coefficient matrix $Df(0)$ as $|x|\to\infty$ if the solution $q(x)$ is localized. Therefore the linearization of such a reaction diffusion stationary problem on $\mathbb{R}^{n}$ about any localized solution $q(x)$ has exponential dichotomies in both the core field and the far field provided that $Df(0)$ has eigenvalues off of the negative real line.
A natural extension would be to perturb the reaction diffusion system with a small inhomogenous nonlinearity,
$0=\Delta u+f(u)+\mu h(x,u,\mu)=\Delta u+Df(q(x))u+\big[f(u)-Df(q(x))u+\mu h(x,u,\mu)\big],$ (22)
and use the core-field and far-field dichotomies for the linearization to construct bounded solutions in both the core field ($r\in(0,1]$) and the far field ($r\in[1,\infty)$). Further analysis from the channel-setting, for example, may be extended to match these solutions and prove the existence of a bounded solution on the whole domain. Of interest would be to characterize how general the inhomogeneity $h$ can be while preserving smooth perturbation of $q(x)$ and absence of essential spectrum.
A similar construction was demonstrated in the channel setting, $(x,y)\in\mathbb{R}\times\Omega$ with $\Omega$ compact in $\mathbb{R}^{n-1}$, by [17] for a spatial evolutionary equation of the form:
@@
where $A$ is an autonomous, linear (differential) operator on $L^{2}(\Omega)$ and $H$ is periodic in $x$. Assuming that a homoclinic solution $q(x)$ exists for $\mu=0$ in (23) and that the linearization about $q$ has an exponential dichotomy on $\mathbb{R}^{+}$ and on $\mathbb{R}^{-}$, [17] constructed a bounded solution on $\mathbb{R}^{+}$ and $\mathbb{R}^{-}$, respectively. Then, by applying a Melnikov argument with Lyapunov-Schmidt techniques, [17] showed that the stable space on $\mathbb{R}^{+}$ and the unstable space on $\mathbb{R}^{-}$ intersect in such a way that a unique bounded solution exists on the whole domain $\mathbb{R}$ when $\mu>0$ is small.
Ongoing and future work aims to extend such an argument to our more general domain with the primary challenge being the higher dimensional kernel that comes from having not just one translational symmetry but at least $n$ symmetries, one for each unbounded direction in state space. With such an argument, the existence of exponential dichotomies in the setting described above would allow us to take previously understood radially symmetric localized solutions and perturb them in non-radially symmetric ways to prove the existence of arbitrary localized stationary patterns.
It would also be of interest to investigate the connection of our approach to the general works on solutions of singularly perturbed nonlinear elliptic problems of the form $\epsilon^{2}\Delta u+V(|x|)u-u|u|^{p-1}=0,$ posed on $\mathbb{R}^{n}$ for suitable $p$, where radial solutions concentrated on spherical subsets related to minima of a certain modified potential have been identified, along with an sequence of non-radial solutions bifurcating as $\epsilon\rightarrow 0^{+}$; see [2, 1] and references therein.
@@
For all $k\geq 0$, $\mu_{k}^{1/2}\leq(1+\mu_{k})^{1/2}$. Likewise, for $r\geq 1$, $\mu_{k}^{1/2}/r\leq(1+\mu_{k})^{1/2}$. Furthermore,
$\sum_{k,j}\mu_{k}\|u_{kj}\|^{2}_{\mathbb{C}^{d}}\leq\|u\|_{H^{1}}^{2}.$
Remark.
Though Lemma A.1 is trivial, since $\mu_{k}\geq 0$ is real valued, we find it convenient to refer to this bound within long calculations for the ease of the reader.